OpenStudy (anonymous):

sqrt(x) = x - 2. solve for x.

OpenStudy (anonymous):

first square the whole thing x=(x-2)^2 x=x^2-4x+4 make equation equal 0 x^2-5x+4=0 use the quadratic formula

OpenStudy (anonymous):

square both sides, solve the resulting quadratic equation, check your answer

OpenStudy (anonymous):

this is a hell of a lot easier than solving partial fractions!!

OpenStudy (anonymous):

haha... this is part of a larger problem

OpenStudy (anonymous):

whew

OpenStudy (anonymous):

i don't think you even need the quadratic formula

OpenStudy (anonymous):

(x-4)(x-1)=0

OpenStudy (anonymous):

\(x^2-5x+4=0\) \((x-4)(x-1)=0\) etc

OpenStudy (anonymous):

of course \(1\) is not a solution, only \(4\) is

OpenStudy (anonymous):

why

OpenStudy (anonymous):

whhhyyy

OpenStudy (anonymous):

@satellite73 why is 1 not a solution

OpenStudy (anonymous):